Convert the given base ten numeral to the indicated base.
4060 to base two

Answer by greenestamps(13200)



Using the base 10 arithmetic you are familiar with, repeatedly divide by 2 and pick off the remainders. Those remainders are the digits of the base 2 representation of the number, starting with the rightmost digit.

4060/2 = 2030 R 0
2030/2 = 1015 R 0
1015/2 = 507 R 1
507/2 = 253 R 1
253/2 = 126 R 1
126/2 = 63 R 0
63/2 = 31 R 1
31/2 = 15 R 1
15/2 = 7 R 1
7/2 = 3 R 1
3/2 = 1 R 1
1/2 = 0 R 1

4060 (base 10) = 111111011100 (Base 2)